OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the system of linear equations below? {2x+y=2 x-4y=10} a- (2,-2) b-(0,1) c-(2 4/7, -3 1/7) d-(-2,6)

OpenStudy (anonymous):

ok solve equation 1 for y \[\Large y=2-2x\] put in the second equation \[\Large x-4(2-2x)=10\] simplify \[\Large x-8+8x=10\] solve this for x then put this value in any of the given equations to find y can you do this ?

OpenStudy (anonymous):

@coolkid2016